Journalist Report Sol 3 – March 14th

After a long day working on our experiments yesterday, we slept like babies. Our nights at the MDRS have generally been quite good, despite the thin mattresses and the out of control heating cycles, which have been pumping hot air into some rooms while others have been left to freeze. Well rested, we woke up in music again (Kanye West’s Mercy this time) and headed downstairs for a work out, this time lead by our biologist Mario who made sure no one missed a single push-up.

Immediately after, four members of our crew embarked on a EVA which took them to the northernmost part of the territory. In this zone, called The Moon, a wide variety of landscapes succeed each other rapidly, going from red striped hills to riverbeds and from sweeping white dunes to areas of sprawling Martian vegetation in just hundreds of meters, which makes for an extremely beautiful (and photogenic) ride. We therefore used this EVA to take more pictures, as well as to collect some samples of these varied soils for Max’s soil testing experiments.

At noon, we had one of the very earth-like, and therefore very delicious meals that Ariane has gotten us accustomed to. This time it was ‘hachis parmentier’, a specialty made with tomatoes, beef, potato mash and grated cheese, cooked in the oven until golden. All of this, of course, was made from freeze-dried foods, but Ariane’s magic touch made it feel right like home.

The rest of the afternoon was dedicated, as usual, to working on our experiments. Mario, with the much welcome help of Max, finished his first hydroponics tower in the greenhab, and planted mint and coriander in it. Sophie managed to find the source of parasites in her muon detector, while Martin dosed the API (active pharmaceutical ingredient) in a drug. On his side, Bastien made a 3D model of the Musk Observatory before embarking on the task of modelising the whole station.

By now, the crew has grown quite accustomed to the daily routine of the MDRS, and things are starting to work smoothly. Despite our confined environment, the mood is still great, and we are organising small events in the evenings, like a karaoke night, game nights, as well as round tables to talk about our experiments and our vision for the project.

See you tomorrow !

Mario

GreenHab Report – March 14th

Crewmember’s name : Mario Sundic (GreenHab officer & journalist)
Date 14/03/2018

Environmental control:

Ambient with door open.
Shade cloth on.
Fan on, without cooling.

Average temperatures: 25°C
Low : 9,5°C
High : 37°C

Hours of supplemental light: 0 hours

Changes to crops: tomatoes are growing and plants are fine and watered

Daily water usage for crops: 10 gal and 10 gal for hydroponics systems

Time(s) of watering for crops: 3

Moringa research observations: ?

Changes to research plants: N/A

I worked on hydroponics systems, completed the first tower, optimized water flow and plants (mint and coriander) are now transplanted on the first tower.

Support/supplies needed: N/A

Operations Report – March 14th

Crew 190 Operations Report 14th March 2018

SOL: 3

Name of person filing report: Bastien BAIX

Non-nominal systems: –

Notes on non-nominal systems: –

Generator (hours run): Turned off at 10:31 am and turned on at 6.47 pm.

Solar – SOC 69 % (Before generator is run at night)

Diesel – ~45%

Propane – 81%

Ethanol Free Gasoline (5 Gallon containers for ATV) – 2 empty jerrycans.

Water (trailer) – 0 gallons.

Water (static) – ~350 gallons

Water (third tank) – 500 gallons

Trailer to Static Pump used – NO

Water (loft) – Static to Loft Pump used – YES

Water Meter: 132 142.7

Toilet tank emptied: YES

ATV’s Used: 2 & 3

Oil Added: NO

ATV Fuel Used: NO

Hours the ATVs were Used today: 2 hours

Notes on ATVs: –

Deimos rover used: NO

Hours: 116.6

Beginning charge: 100%

Ending charge: –

Currently charging: NO

Notes on Deimos rover: flat tire, not yet repared

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: NO

Hours: 24.8

Beginning charge: 100%

Ending charge: –

Currently charging: NO

Opportunity rover used: YES

Hours: 21.5

Beginning charge: 37%

Ending charge: 100%

Currently charging: YES

Curiosity rover used: NO

Hours: 17,2

Beginning charge: 100%

Ending charge: –

Currently charging: YES

HabCar used and why, where: –

General notes and comments: –

Summary of internet: ~ 300 Mb remaining

Summary of suits and radios: –

Summary of Hab operations: Ariane & Bastien working on their experiments

Summary of GreenHab operations: Mario (+Maximilien) & Fred working on their experiments

Summary of ScienceDome operations: Martin, Sophie & Michael working on their experiments.

Summary of RAMM operations: –

Summary of health and safety issues: a crewmember had a headache and another one cut himself/herself

Questions, concerns and requests to Mission Support: –

Crew 190 Sol Summary 14Mar2018

[title Sol Summary -March 14th]

Crew 190 Sol Summary Report
14March2018
Sol 3

Summary Title:
Exploration X Science X Pancakes

Michael Saint-Guillain (XO)

Mission Status:
Ready to continue simulation on Sol 4, 08:00

Sol Activity Summary:
8:00 Gym
8:20 Breakfast
9:10 EVA: Sophie, Mario, Frederic, Michael
9:15 Cleaning and social activity preparation: Maximilien, Martin, Ariane, Bastien
10:55 EVA ended; Crew meeting before lunch
13:00 Lunch
14:30 Scientific work: muongraphy (Sophie), video extraction & compilation (Bastien), chemistry/biology (Ariane, Martin, Maximilien, Frederic), GreenHab (Mario), operations research and academic work (Michael)
18:00 Redaction of the reports
19:00 CapCom

Optimized Look Ahead Plan:
08:00 – 08:20 Gym
08:20 – 09:00 Breakfast
09:30 – 12:30 EVA: Ariane, Bastien, Maximilien, Michael
09:30 – 12:30 Scientific work: Martin (pharmacology), Sophie (muography, telescope), Mario & Frederic (green hab)
12:30 – 14:00 Lunch
14:00 – 18:00 Back to scientific work: Ariane & Maximilien (chemistry), Bastien (cartography), Michael (operations research)
18:00 – 19:00 Redaction of the reports
19:00 Capcom

Forecast:
To be updated March 15 2018

Anomalies in work:
None

Weather:
Cloudy and a bit cold in the morning. Showers in the afternoon, preventing Sophie from telescope activity.

Crew Physical Status:
In increasingly good shape (thanks to the Gym!) and enthusiasm (pancakes).

EVA:
See EVA report. Mainly soil sampling and pictures.

Reports to be file:
Commander report
Journalist report
EVA#4 report
EVA#5 request (Sol4)
Engineer report
GreenHab
HSO report
Astronomy report

Support Requested:
None

Sol 3 Commander Report – March 14th

Crew 190 – sol3 – crew commander report 14-03-2018

As a habit for 3 days, we started our day by a gym session given by Mario. Push up, abs and more to stay fit during the whole mission! After our exercise session and a pancake time, four of us left for an EVA. Unfortunately, they forgot to take some parts of the material before going in the air lock. Therefore, we had to do the outing procedure once again after taking the required material. Frédéric, Michael, Mario and Sophie left with their material and two flags. A Belgian one and the second one with the name and logo of our Belgian university!

After riding for a while with the rovers, they arrived to “the moon” area where they took some soil samples for me and pictures with flags. When they were back to the Hab, Ariane and Michael started to cook the lunch for the whole crew. Afterwards, everyone went to his scientific work for the entire afternoon. Michael did academic work and operational scheduling for the crew. He is still calculating the schedule of tomorrow and it will be soon available. Regarding the part of the team working in the science dome, Martin got good results for his molecules analysis and Ariane was able to follow her sourdough fermentation using pH measurement. Sophie tried to use her muon detector into the science dome but the noise produced by other electrical equipment disturbed her measurement. She will try to perform a measurement away from the station to avoid perturbation coming from equipment.

Regarding the green hab, Mario has finished his first hydroponic tower. Some problem of water leaks appear but we found a solution to fix it and make the hydroponic tower effective. Now the station enjoy a new efficient possibility to make crops growing!

Regarding the mood in the team, it is still good and no problem appeared until now. Nobody complain of our confined environment and wish to go back earth soon!

For tonight, we planned a personal story session to tell the others what we were doing before being sent as the first humans to Mars.

Tomorrow, another EVA will be planned with Ariane as EVA leader (see EVA request). We will continue to work on our scientific projects and find new data about Mars!

EVA#4 Report – March 14th

EVA #4 – 14th March 2018 – Sol 3

Crew members: Sophie Wuyckens (EVA leader), Michael Saint-Guillain, Frédéric Peyrusson, Mario Sundic

Objectives:

– Soil samples harvest by Fred for Maximilien’s experiment.

– Hill localization for Sophie’s experiment.

– Pictures

Operations:

9:01 – Entering the airlock

9:08 – Leaving the MDRS with 2 ATVs and 1 Rover (Opportunity)

9:26 – Arrival at Beige Moon for pictures time

9:43 – Second stop just before splitted roads again for picture time

10:02 – Third stop for soil samples harvest by Fred for Maximilien’s experiment near Widow’s Peak

10:19 – Fourth stop for pictures time with flags

10:25 – Back to the station

10:50 – Arrival at the Hab

10:55– Entering the airlock

10:58– Mission finished

Narrative:

Our fourth EVA took place on soil 3 at 9:00. I am Sophie Wuyckens and I wanted to be part of this EVA in order to find a convenient hill where I could operate my detector.

The EVA has not started very well. Indeed, we were a bit tired and forgot everything ; the belgian flag, our university flag, Mario’s gloves, a notepad (to write operation hours). This is why we stayed a bit longer in the airlock compared to the last times. But finally we made it to outside and started to put firstly our sponsors posters on the rover and ATVs in order to get great pictures for our sponsors. Then we follow the main road and then White Moon direction and stop at Beige Moon. There Mario took pictures of us. I kept an eye on hills that we passed in front of with our vehicles. Unfortunately, we lost a sponsor poster. We continued our way and just before the connecting roads, we made a second stop because the landscape was amazing. Near Widow’s Peak, we stopped a third time because Fred wanted to harvest some soil samples for Maximilien’s experiment. We continued our way along Snake River. Finally, we made a last stop near Toothy Ridge to take last pictures with the flags. After that, we decided to return to the station. The return was very calm. We have never found our poster whereas we have looked for it for a while. Hoping next mission will find it. With regards to my experiment, I found some interesting places near Gray Moon. I look forward to operate the detector over there.

The EVA finished just on time. We hadn’t walk much but it was a good and efficient EVA.

Vehicles

– Opportunity rover

– ATVs 2&3